leilanithomas2024:

v + 4 - 2 = -10 + 6v + 6 + 6

mxddi3:

first you need to combine like terms on both sides

leilanithomas2024:

on both sides or i have to choose one?

mxddi3:

\[v+4-2=-10+6v+6+6\] on the left combine the numbers and on the right combine the numbers

leilanithomas2024:

so on the left is it 2v?

leilanithomas2024:

and the right 28?

leilanithomas2024:

v^

mxddi3:

no. you do not combine the 6v with the numbers because they have no variable. on the left, you combine 4-2 and on the right you combine -10+6+6

leilanithomas2024:

ok 2 and 2?

mxddi3:

so on the left you will have : \[v+2\] and on the right you will have: \[6v+2\] so now solve: \[v+2=6v+2\]

leilanithomas2024:

oh its 0

leilanithomas2024:

v=0

mxddi3:

correct!

leilanithomas2024:

thank you

mxddi3:

yw!